What is the value of the expression? tan (-5pi/6) I'm not sure if I wrote that correctly.
Answer by lwsshak3(11628)

What is the value of the expression? tan (-5pi/6)
Starting from zero on the unit circle, rotate clockwise (5/6)π
You will end up with a reference angle of π/6 in quadrant III where tan>0
tan(-5π/6)=tan(π/6)=√3/3